FTIR ISOTOPIC AND DFT STUDIES OF SiC_{5} TRAPPED IN SOLID Ar

\underline{T. H. LE}, and W. R. M. GRAHAM, Molecular Physics Laboratory, Department of Physics and Astronomy, Texas Christian University, Fort Worth, TX 76129.

We report here the first results from a Fourier transform infrared (FTIR) and density functional theory (DFT) study on SiC_{5}. This species has been produced by ablating a rod made of silicon and graphite powder with a Nd:YAG laser and trapping the products in solid Ar at ∼ 15 K. In the only previous observation of SiC_{5}, McCarthy et al. reported the microwave rotational spectrum^{a}. In the present work, extensive FTIR measurements of vibrational frequencies and isotopic shifts from ^{13}C enrichment and the naturally occurring ^{29,30}Si isotopes have been compared with the predictions of density functional theory calculations at the B3LYP/cc-pVDZ level. The excellent agreement between experiment and theory has enabled the assignment of the ν_{4}(σ) fundamental of linear SiC_{5} at 936.9 ± 0.2 cm^{−1}. This information may help in identifying SiC_{5} in circumstellar and interstellar environments.

aM.C. McCarthy, A.J. Apponi, C.A. Gottlieb, and P. Thaddeus, Astrophys. J. \underline{538}, 766 (2000)